A projectile is launched straight up with an initial volocity of 96 ft per sec. How many seconds to reach 144 feet? 2.5 sec, 3 sec, 3.5 sec, or 4 sec??
Thanks

Answer by stanbon(75887)

A projectile is launched straight up with an initial
velocity of 96 ft per sec.
How many seconds to reach 144 feet?
------------
h(t) = -16t^2+96t
---
144 = -16t^2+96t
16t^2-96t-144 = 0
-----
t^2-6t-9 = 0
(t-3)^3 = 0
time = 3 seconds.
-----------------------
Cheers,
Stan H.
==============
2.5 sec, 3 sec, 3.5 sec, or 4 sec??